# A Comprehensive Survey of Website Fingerprinting Attacks and Defenses in Tor: Advances and Open Challenges

### v1

@@ III-B1 Traffic Timing
Timing features capture temporal patterns in network traffic that can reveal identifying characteristics of a website, even when the content is encrypted. As shown in Figure 4, one commonly used metric is packet inter-arrival time [37], which measures the time interval $\Delta t$ between consecutive packets or groups of packets. Similarly, the time difference between traffic traces and the round-trip time (RTT) provides insights into latency and responsiveness in the communication path [62, 65]. The time gap [62] reflects idle periods within a trace, often corresponding to content loading phases or user interactions. Other measures, such as the ratio of transmission time to total trace time [58, 9, 63, 64], summarize the distribution of active and idle periods within a session. Overall, these timing features help adversaries or analysts distinguish between websites by exploiting subtle but consistent timing signatures.
Timing features present a challenging surface to protect. Unlike packet sizes or directions, which can be easily padded or reshaped, time-based features are inherently tied to user interaction, network conditions, and server response behaviors. Defenses such as adaptive padding [77, 16], timing obfuscation [53, 72], and randomized delays [78] have been proposed to mitigate timing-based attacks. However, these approaches often introduce latency or degrade performance. As a result, incorporating traffic timing awareness in both attacks and defenses is now essential for evaluating the real-world feasibility and robustness of WF techniques.
@@ III-B2 Traffic Length
Traffic length features capture the overall size or volume of a traffic trace, typically measured in terms of the number of packets, cells, or bytes exchanged during a browsing session [9, 36]. In WF attacks, traffic length is often highly informative because different websites require varying amounts of data to load. For instance, a media-rich website generally produces longer traces than a minimalist blog. Adversaries exploit this variability by treating traffic length as a coarse but effective discriminative feature, particularly in closed-world settings where the set of candidate websites is fixed and well-known. Even under encryption or anonymity systems, the total session length can still reveal strong clues about the visited website. However, in real-world scenarios, attackers face challenges in identifying the precise start and end points of a webpage load. To address this, some studies [64, 65, 62] approximate traffic length using time gap-based measurements. As shown in Figure 4, $T/L$ represents the fixed time gap or traffic length for captured Tor traffic flow. While this approach helps mitigate alignment issues, it can reduce accuracy when two different webpages produce traffic traces with similar lengths and time gaps.
@@ III-B3 Cell Size
In Tor, traffic is encapsulated into fixed-size units called cells (typically 512 bytes), so attackers cannot directly infer information from individual cells [37]. In early studies, researchers captured Tor traffic in a closed-world setting, where the total cell count and packet length for each webpage were known. Thus, cell size is widely used in traditional ML-based WF attacks [9, 7, 35, 79, 80]. However, in real-world scenarios, attackers cannot precisely measure the total traffic length because the start and end Tor cells are unknown. Consequently, recent website fingerprinting studies seldom rely on cell size for website classification.
@@ III-B4 Packet Ordering
Packet ordering features refer to the sequence in which packets or cells are transmitted and received during a browsing session, capturing the precise arrangement of request-response interactions between a client and a server [7, 35]. In WF attacks, attackers can analyze these ordered sequences to uncover consistent patterns that are unique to specific websites. For instance, certain pages may always begin with a small outgoing request followed by a predictable series of large incoming responses, or they may load specific resources in a particular order. By modeling the order of packets, along with their direction and timing, ML algorithms can effectively fingerprint websites from encrypted Traffic. This makes packet ordering a subtle yet powerful indicator of web activity, especially when used in conjunction with other features, such as inter-packet timing and burst behavior [35, 64].
From a defensive standpoint, disrupting recognizable packet ordering is a complex task, as it requires altering the natural flow of web traffic without degrading usability or breaking functionality. Defenses that aim to obscure packet order often rely on techniques such as injecting dummy packets [55, 67, 68], randomized packet delays [58, 69], or traffic shaping [22, 23, 19], which attempt to introduce noise into the observed sequence of packets. However, these approaches come with significant trade-offs, including added latency, reduced performance, and the risk of interfering with application-layer protocols. Moreover, attackers can sometimes adapt to these defenses by leveraging probabilistic models that tolerate small variations in ordering. As a result, while packet ordering is a less frequently discussed feature compared to direction or timing, it plays a critical role in advanced WF attacks and must be carefully considered in the design of robust and practical defenses.
@@ III-B5 Traffic Bursts
Traffic bursts are defined as sequences of consecutive packets in the same direction, either incoming or outgoing, grouped between directional switches. In WF attacks, burst patterns capture higher-level structural information about how websites load resources, reflecting how browsers and servers exchange data in grouped transactions. For example, a user request (outgoing burst) might be followed by multiple server responses (incoming burst), forming a distinct burst sequence unique to the website [71]. These burst patterns, characterized by burst size, burst count, and burst order, are often stable across visits and provide a rich feature set for ML and DL models. SOTA attacks leverage bursts as highly discriminative features for classifying encrypted traces, prompting most WF defenses to focus on obscuring them.
Defending against burst-based fingerprinting requires strategies that disrupt the predictability of burst structure without significantly impacting user experience. Techniques like burst molding, dummy burst insertion, or splitting and merging of bursts have been proposed to obscure these patterns. Defenses such as Walkie-Talkie [45] and WTF-PAD [16] specifically target burst behavior by reshaping traffic to mimic the burst patterns of multiple websites or injecting randomness into burst timing and size. However, these approaches often introduce challenges such as increased latency, bandwidth overhead, or implementation complexity. Additionally, attackers demonstrated that even partially obscured burst patterns can leak identifiable information when combined with other features. Therefore, traffic bursts represent a critical feature class in both WF attacks and defenses, requiring sophisticated and adaptive countermeasures to mitigate their fingerprinting potential effectively.
@@ III-B6 Traffic Direction
Traffic direction features refer to the sequence and pattern of incoming and outgoing packets or cells during a user’s interaction with a website. In WF attacks, the direction of traffic, represented as sequences of +1 (outgoing) and -1 (incoming), forms one of the most distinguishing features for identifying websites [12, 39, 40, 33].
@@ III-B7 Coarse-grained Features
Coarse-grained features in website fingerprinting refer to high-level traffic statistics that enable website identification without requiring decryption. They exploit differences in website loading behaviors, producing unique traffic patterns that persist even under encryption and fine-grained defenses. Per-direction bandwidth reveals the ratio of incoming to outgoing data. For example, a video streaming site will have a massive incoming bandwidth and a relatively small outgoing bandwidth, while a file-upload site will exhibit the opposite. Traffic burstiness refers to the tendency of data transfer to occur in bursts rather than at a constant rate. Even if some defense models try to obfuscate burst patterns in Tor traffic, they often fail to address traffic burstiness and burst bandwidth. These coarse-grained features can still leak significant information and allow for website fingerprinting [58].
@@ III-C WF Attacks
Website fingerprinting (WF) [11, 35, 36, 27, 81], also known as a WF attack, is a type of traffic analysis technique used to infer which websites a user is visiting, even when the traffic is encrypted and anonymized through tools like VPNs or the Tor network. We compare the current representative encrypted traffic protocols in TABLE II. By analyzing patterns in network metadata, such as packet sizes, timing, and sequences, attackers can bypass encryption (e.g., HTTPS, VPNs, or Tor) and infer browsing activity. These attacks pose significant risks to user anonymity, particularly in environments with heavy surveillance or censorship. Therefore, WF attacks pose a significant threat to user anonymity on the Tor network. Adversaries can exploit metadata patterns, such as packet sizes, timing, and sequences, to infer the websites a user visits. These attacks undermine Tor’s privacy guarantees, particularly for users in censored or surveilled environments.
@@ III-C1 Traditional Machine Learning Methods
Wagner et al. [82] designed and successfully implemented the first website fingerprinting and conducted preliminary experiments on a small number of websites. The concept of the WF attack was first proposed by Hintz [83] in 2002. They used web object size as fingerprints and verified the validity of WF attacks on the encrypted web proxy SafeWeb in small-scale experiments. Yan and Kaur [14] identify visited websites by analyzing TCP/IP header data, even when content is encrypted. The study focuses on a limited set of features by conducting an exhaustive analysis of TCP/IP features across diverse communication scenarios. The goal is to understand the full extent of website fingerprintability and identify previously unknown informative features. The early WF attacks are mainly performed against HTTPS [11], encrypted web proxies [84, 58], OpenSSH [85, 66, 86, 87], OpenVPN [86]. In the case of HTTPS, OpenSSH, and OpenVPN, researchers have gathered TCP/IP data packets generated when visiting HTTPS websites. The traffic is then transformed into a vector based on the order and direction of the data packets, as shown in Figure LABEL:subfig:tcptraffic. Until 2009, Herrmann et al. [8] performed the attack on the Tor network for the first time, which was based on a Naïve Bayes classifier with the features from the normalized frequency distribution of packet size. However, only 2.96% accuracy is obtained. Since then, WF attacks against Tor have become a research hotspot. And machine learning was applied to improve accuracy, including support vector machines (SVMs) [4, 7, 35], hidden markov models (HMM) [7, 35], and k-nearest neighbors (k-NN) [9]. We briefly introduce traditional ML methods as follows.
@@
  • Naïve Bayes [88] is a supervised ML algorithm based on Bayes’ theorem and the assumption of conditional independence among features. It calculates the posterior probability of each class given the observed features and selects the class with the highest probability, making it suitable for both binary and multiclass classification tasks. Due to its simplicity, scalability, and effectiveness in handling high-dimensional data, Naïve Bayes is widely used as a baseline model in ML.
@@
  • SVMs [89] are a class of supervised learning algorithms widely used for classification and regression tasks in ML. It works by finding the optimal hyperplane that best separates data points of different classes with the maximum margin. For non-linearly separable data, SVM uses kernel functions, such as polynomial, Gaussian, and sigmoid, to transform the data into a higher-dimensional space where separation is possible. SVM is known for its effectiveness in high-dimensional spaces and its ability to generalize well with limited training data.
@@
  • k-NN [90] is a commonly used supervised learning method. It is a non-parametric, instance-based algorithm that makes no assumptions about the data distribution and uses the training data directly for prediction. Its simplicity and effectiveness allow it to identify the k closest samples in the training data and make predictions based on the majority label (for classification) or the average value (for regression) of these neighbors.
@@
  • Decision Tree [91] is another supervised learning algorithm used for both classification and regression tasks. It models decisions and their possible consequences in a tree-like structure, where each internal node represents a test on a feature, each branch corresponds to the outcome of the test, and each leaf node represents a final prediction or decision. ID3, C4.5, and CART are commonly used algorithms for constructing decision trees.
@@
  • Random Forest [92] builds upon the concept of decision trees by creating a “forest” of multiple trees. Random Forest achieves improved accuracy and robustness by aggregating the predictions from many diverse trees. This approach helps overcome the common problem of overfitting that individual decision trees often face.
@@ III-C2 Deep Learning Methods
Recently, a new class of WF attacks leveraging deep learning (DL) has emerged, aiming to enhance the accuracy, effectiveness, and robustness of Tor traffic analysis [10, 93, 94, 95, 96, 12, 97, 98, 87]. These DL-based approaches have consistently achieved accuracies exceeding 90%, establishing themselves as SOTA WF attacks and serving as benchmarks for evaluating subsequent attacks and defenses. By exploiting structured input features and harnessing the powerful representation learning capabilities of DL architectures, these methods can automatically extract complex patterns from encrypted traffic, leading to superior classification performance. A brief overview of the representative DL models used in WF attacks is presented below.
@@
  • Convolutional Neural Networks (CNNs) [99] are a class of DL models particularly effective for automatically extracting hierarchical spatial and temporal patterns without requiring handcrafted features. They use convolutional layers to automatically learn and extract local features by applying small filters across the input, capturing spatial or temporal patterns. Pooling layers then reduce dimensionality while preserving essential information, and fully connected layers integrate the extracted features for classification or regression tasks. CNNs excel at detecting hierarchical patterns, making them widely used in fields such as computer vision, natural language processing, and traffic analysis.
@@
  • Long Short-Term Memory (LSTM) [100] is a type of recurrent neural network (RNN) designed to capture long-range dependencies in sequential data effectively. By incorporating memory cells and gating mechanisms (input, output, and forget gates), LSTMs can selectively retain or discard information over time, effectively mitigating the vanishing and exploding gradient problems common in traditional RNNs. This capability makes them particularly effective for modeling temporal patterns and context in tasks such as natural language processing, speech recognition, and traffic analysis.
@@
  • Generative Adversarial Networks (GANs) [101] are a class of DL models consisting of a generator and a discriminator that are trained simultaneously in a competitive process. The generator creates synthetic data intended to resemble real data, while the discriminator attempts to distinguish between real and generated samples. Through this adversarial training, the generator learns to produce highly realistic outputs. GANs are widely used for tasks such as image synthesis, data augmentation, and anomaly detection.
@@
  • Transformers [102] are a type of DL architecture designed to handle sequential data by relying entirely on an attention mechanism, rather than recurrence or convolution. The attention mechanism enables the model to weigh the importance of different elements in a sequence, capturing long-range dependencies efficiently. Transformers are highly parallelizable and have achieved SOTA results in natural language processing, computer vision, and other sequence modeling tasks.
@@
  • Graph Neural Networks (GNNs) [103] are a class of neural networks designed to work directly with graph-structured data. They learn to represent nodes, edges, and entire graphs by aggregating and transforming information from a node’s neighbors. This enables GNNs to capture complex relationships and patterns in data where connections are crucial, such as social networks, molecules, and transportation systems. GNNs have been successfully applied in various domains, including recommendation systems, chemistry, and network analysis.

### v2

@@ III-B1 Traffic Timing
Timing features capture temporal patterns in network traffic that can reveal identifying characteristics of a website, even when the content is encrypted. As shown in Figure 4, one commonly used metric is packet inter-arrival time [wang2016realistically], which measures the time interval $\Delta t$ between consecutive packets or groups of packets. Similarly, the time difference between traffic traces and the round-trip time (RTT) provides insights into latency and responsiveness in the communication path [xu2018multi, yin2021automated]. The time gap [xu2018multi] reflects idle periods within a trace, often corresponding to content loading phases or user interactions. Other measures, such as the ratio of transmission time to total trace time [dyer2012peek, wang2014effective, kwon2015circuit, attarian2019adawfpa], summarize the distribution of active and idle periods within a session. Overall, these timing features help adversaries or analysts distinguish between websites by exploiting subtle but consistent timing signatures.
Timing features present a challenging surface to protect. Unlike packet sizes or directions, which can be easily padded or reshaped, time-based features are inherently tied to user interaction, network conditions, and server response behaviors. Defenses such as adaptive padding [witwer2022padding, juarez2016toward], timing obfuscation [shmatikov2006timing, lu2023lightweight], and randomized delays [hong2022website] have been proposed to mitigate timing-based attacks. However, these approaches often introduce latency or degrade performance. As a result, incorporating traffic timing awareness in both attacks and defenses is now essential for evaluating the real-world feasibility and robustness of WF techniques.
@@ III-B2 Traffic Length
Traffic length features capture the overall size or volume of a traffic trace, typically measured in terms of the number of packets, cells, or bytes exchanged during a browsing session [wang2014effective, hayes2016k]. In WF attacks, traffic length is often highly informative because different websites require varying amounts of data to load. For instance, a media-rich website generally produces longer traces than a minimalist blog. Adversaries exploit this variability by treating traffic length as a coarse but effective discriminative feature, particularly in closed-world settings where the set of candidate websites is fixed and well-known. Even under encryption or anonymity systems, the total session length can still reveal strong clues about the visited website. However, in real-world scenarios, attackers face challenges in identifying the precise start and end points of a webpage load. To address this, some studies [attarian2019adawfpa, yin2021automated, xu2018multi] approximate traffic length using time gap-based measurements. As shown in Figure 4, $T/L$ represents the fixed time gap or traffic length for captured Tor traffic flow. While this approach helps mitigate alignment issues, it can reduce accuracy when two different webpages produce traffic traces with similar lengths and time gaps.
@@ III-B3 Cell Size
In Tor, traffic is encapsulated into fixed-size units called cells (typically 512 or 514 bytes), so attackers cannot directly infer information from individual cells [wang2016realistically]. In early studies, researchers captured Tor traffic in a closed-world setting, where the total cell count and packet length for each webpage were known. Thus, cell size is widely used in traditional ML-based WF attacks [wang2014effective, cai2012touching, wang2013improved, zhuo2017website, cui2019revisiting]. However, in real-world scenarios, attackers cannot precisely measure the total traffic length because the start and end Tor cells are unknown. Consequently, recent website fingerprinting studies seldom rely on cell size for website classification.
@@ III-B4 Packet Ordering
Packet ordering features refer to the sequence in which packets or cells are transmitted and received during a browsing session, capturing the precise arrangement of request-response interactions between a client and a server [cai2012touching, wang2013improved]. In WF attacks, attackers can analyze these ordered sequences to uncover consistent patterns that are unique to specific websites. For instance, certain pages may always begin with a small outgoing request followed by a predictable series of large incoming responses, or they may load specific resources in a particular order. By modeling the order of packets, along with their direction and timing, ML algorithms can effectively fingerprint websites from encrypted Traffic. This makes packet ordering a subtle yet powerful indicator of web activity, especially when used in conjunction with other features, such as inter-packet timing and burst behavior [wang2013improved, attarian2019adawfpa].
From a defensive standpoint, disrupting recognizable packet ordering is a complex task, as it requires altering the natural flow of web traffic without degrading usability or breaking functionality. Defenses that aim to obscure packet order often rely on techniques such as injecting dummy packets [pulls2020towards, abusnaina2020dfd, kadianakis2021tor], randomized packet delays [dyer2012peek, cai2014cs], or traffic shaping [chan2018website, al2019bimorphing, shen2024real], which attempt to introduce noise into the observed sequence of packets. However, these approaches come with significant trade-offs, including added latency, reduced performance, and the risk of interfering with application-layer protocols. Moreover, attackers can sometimes adapt to these defenses by leveraging probabilistic models that tolerate small variations in ordering. As a result, while packet ordering is a less frequently discussed feature compared to direction or timing, it plays a critical role in advanced WF attacks and must be carefully considered in the design of robust and practical defenses.
@@ III-B5 Traffic Bursts
Traffic bursts are defined as sequences of consecutive packets in the same direction, either incoming or outgoing, grouped between directional switches. In WF attacks, burst patterns capture higher-level structural information about how websites load resources, reflecting how browsers and servers exchange data in grouped transactions. For example, a user request (outgoing burst) might be followed by multiple server responses (incoming burst), forming a distinct burst sequence unique to the website [bahramali2023realistic]. These burst patterns, characterized by burst size, burst count, and burst order, are often stable across visits and provide a rich feature set for ML and DL models. SOTA attacks leverage bursts as highly discriminative features for classifying encrypted traces, prompting most WF defenses to focus on obscuring them.
Defending against burst-based fingerprinting requires strategies that disrupt the predictability of burst structure without significantly impacting user experience. Techniques like burst molding, dummy burst insertion, or splitting and merging of bursts have been proposed to obscure these patterns. Defenses such as Walkie-Talkie [wang2017walkie] and WTF-PAD [juarez2016toward] specifically target burst behavior by reshaping traffic to mimic the burst patterns of multiple websites or injecting randomness into burst timing and size. However, these approaches often introduce challenges such as increased latency, bandwidth overhead, or implementation complexity. Additionally, attackers demonstrated that even partially obscured burst patterns can leak identifiable information when combined with other features. Therefore, traffic bursts represent a critical feature class in both WF attacks and defenses, requiring sophisticated and adaptive countermeasures to mitigate their fingerprinting potential effectively.
@@ III-B6 Traffic Direction
Traffic direction features refer to the sequence and pattern of incoming and outgoing packets or cells during a user’s interaction with a website. In WF attacks, the direction of traffic, represented as sequences of +1 (outgoing) and -1 (incoming), forms one of the most distinguishing features for identifying websites [rahman2019tik, shi2025multiscale, zou2024relation, zhao2024towards].
@@ III-B7 Coarse-grained Features
Coarse-grained features in website fingerprinting refer to high-level traffic statistics that enable website identification without requiring decryption. They exploit differences in website loading behaviors, producing unique traffic patterns that persist even under encryption and fine-grained defenses. Per-direction bandwidth reveals the ratio of incoming to outgoing data. For example, a video streaming site will have a massive incoming bandwidth and a relatively small outgoing bandwidth, while a file-upload site will exhibit the opposite. Traffic burstiness refers to the tendency of data transfer to occur in bursts rather than at a constant rate. Even if some defense models try to obfuscate burst patterns in Tor traffic, they often fail to address traffic burstiness and burst bandwidth. These coarse-grained features can still leak significant information and allow for website fingerprinting [dyer2012peek].
@@ III-C WF Attacks
Website fingerprinting (WF) [panchenko2016website, wang2013improved, hayes2016k, aminuddin2023rise, fu2024detecting], also known as a WF attack, is a type of traffic analysis technique used to infer which websites a user is visiting, even when the traffic is encrypted and anonymized through tools like VPNs or the Tor network. We compare the current representative encrypted traffic protocols in TABLE II. By analyzing patterns in network metadata, such as packet sizes, timing, and sequences, attackers can bypass encryption (e.g., HTTPS, VPNs, or Tor) and infer browsing activity. These attacks pose significant risks to user anonymity, particularly in environments with heavy surveillance or censorship. Therefore, WF attacks pose a significant threat to user anonymity on the Tor network. Adversaries can exploit metadata patterns, such as packet sizes, timing, and sequences, to infer the websites a user visits. These attacks undermine Tor’s privacy guarantees, particularly for users in censored or surveilled environments.
@@ III-C1 Traditional Machine Learning Methods
Wagner et al. [wagner1996analysis] designed and successfully implemented the first website fingerprinting and conducted preliminary experiments on a small number of websites. The concept of the WF attack was first proposed by Hintz [hintz2002fingerprinting] in 2002. They used web object size as fingerprints and verified the validity of WF attacks on the encrypted web proxy SafeWeb in small-scale experiments. Yan and Kaur [yan2018feature] identify visited websites by analyzing TCP/IP header data, even when content is encrypted. The study focuses on a limited set of features by conducting an exhaustive analysis of TCP/IP features across diverse communication scenarios. The goal is to understand the full extent of website fingerprintability and identify previously unknown informative features. The early WF attacks are mainly performed against HTTPS [panchenko2016website], encrypted web proxies [sun2002statistical, dyer2012peek], OpenSSH [bissias2005privacy, liberatore2006inferring, lu2010website, guo2021website], OpenVPN [lu2010website]. In the case of HTTPS, OpenSSH, and OpenVPN, researchers have gathered TCP/IP data packets generated when visiting HTTPS websites. The traffic is then transformed into a vector based on the order and direction of the data packets, as shown in Figure LABEL:subfig:tcptraffic. Until 2009, Herrmann et al. [herrmann2009website] performed the attack on the Tor network for the first time, which was based on a Naïve Bayes classifier with the features from the normalized frequency distribution of packet size. However, only 2.96% accuracy is obtained. Since then, WF attacks against Tor have become a research hotspot. And machine learning was applied to improve accuracy, including support vector machines (SVMs) [panchenko2011website, cai2012touching, wang2013improved], hidden markov models (HMM) [cai2012touching, wang2013improved], and k-nearest neighbors (k-NN) [wang2014effective]. We briefly introduce traditional ML methods as follows.
@@
  • Naïve Bayes [rish2001empirical] is a supervised ML algorithm based on Bayes’ theorem and the assumption of conditional independence among features. It calculates the posterior probability of each class given the observed features and selects the class with the highest probability, making it suitable for both binary and multiclass classification tasks. Due to its simplicity, scalability, and effectiveness in handling high-dimensional data, Naïve Bayes is widely used as a baseline model in ML.
@@
  • SVMs [pisner2020support] are a class of supervised learning algorithms widely used for classification and regression tasks in ML. It works by finding the optimal hyperplane that best separates data points of different classes with the maximum margin. For non-linearly separable data, SVM uses kernel functions, such as polynomial, Gaussian, and sigmoid, to transform the data into a higher-dimensional space where separation is possible. SVM is known for its effectiveness in high-dimensional spaces and its ability to generalize well with limited training data.
@@
  • k-NN [zhang2016introduction] is a commonly used supervised learning method. It is a non-parametric, instance-based algorithm that makes no assumptions about the data distribution and uses the training data directly for prediction. Its simplicity and effectiveness allow it to identify the k closest samples in the training data and make predictions based on the majority label (for classification) or the average value (for regression) of these neighbors.
@@
  • Decision Tree [myles2004introduction] is another supervised learning algorithm used for both classification and regression tasks. It models decisions and their possible consequences in a tree-like structure, where each internal node represents a test on a feature, each branch corresponds to the outcome of the test, and each leaf node represents a final prediction or decision. ID3, C4.5, and CART are commonly used algorithms for constructing decision trees.
@@
  • Random Forest [biau2016random] builds upon the concept of decision trees by creating a “forest” of multiple trees. Random Forest achieves improved accuracy and robustness by aggregating the predictions from many diverse trees. This approach helps overcome the common problem of overfitting that individual decision trees often face.
@@ III-C2 Deep Learning Methods
Recently, a new class of WF attacks leveraging deep learning (DL) has emerged, aiming to enhance the accuracy, effectiveness, and robustness of Tor traffic analysis [abe2016fingerprinting, sirinam2018deep, rimmer2017automated, sirinam2019triplet, bhat2018var, rahman2019tik, chen2021few, guo2021deep, guo2021website]. These DL-based approaches have consistently achieved accuracies exceeding 90%, establishing themselves as SOTA WF attacks and serving as benchmarks for evaluating subsequent attacks and defenses. By exploiting structured input features and harnessing the powerful representation learning capabilities of DL architectures, these methods can automatically extract complex patterns from encrypted traffic, leading to superior classification performance. A brief overview of the representative DL models used in WF attacks is presented below.
@@
  • Convolutional Neural Networks (CNNs) [lecun2002gradient] are a class of DL models particularly effective for automatically extracting hierarchical spatial and temporal patterns without requiring handcrafted features. They use convolutional layers to automatically learn and extract local features by applying small filters across the input, capturing spatial or temporal patterns. Pooling layers then reduce dimensionality while preserving essential information, and fully connected layers integrate the extracted features for classification or regression tasks. CNNs excel at detecting hierarchical patterns, making them widely used in fields such as computer vision, natural language processing, and traffic analysis.
@@
  • Long Short-Term Memory (LSTM) [graves2012long] is a type of recurrent neural network (RNN) designed to capture long-range dependencies in sequential data effectively. By incorporating memory cells and gating mechanisms (input, output, and forget gates), LSTMs can selectively retain or discard information over time, effectively mitigating the vanishing and exploding gradient problems common in traditional RNNs. This capability makes them particularly effective for modeling temporal patterns and context in tasks such as natural language processing, speech recognition, and traffic analysis.
@@
  • Generative Adversarial Networks (GANs) [goodfellow2020generative] are a class of DL models consisting of a generator and a discriminator that are trained simultaneously in a competitive process. The generator creates synthetic data intended to resemble real data, while the discriminator attempts to distinguish between real and generated samples. Through this adversarial training, the generator learns to produce highly realistic outputs. GANs are widely used for tasks such as image synthesis, data augmentation, and anomaly detection.
@@
  • Transformers [khan2022transformers] are a type of DL architecture designed to handle sequential data by relying entirely on an attention mechanism, rather than recurrence or convolution. The attention mechanism enables the model to weigh the importance of different elements in a sequence, capturing long-range dependencies efficiently. Transformers are highly parallelizable and have achieved SOTA results in natural language processing, computer vision, and other sequence modeling tasks.
@@
  • Graph Neural Networks (GNNs) [scarselli2008graph] are a class of neural networks designed to work directly with graph-structured data. They learn to represent nodes, edges, and entire graphs by aggregating and transforming information from a node’s neighbors. This enables GNNs to capture complex relationships and patterns in data where connections are crucial, such as social networks, molecules, and transportation systems. GNNs have been successfully applied in various domains, including recommendation systems, chemistry, and network analysis.